1000 euros is invested at 3% per annum interest, compounded monthly. Calculate the minimum number of months required for the value to exceed 1300 euros.
I know that this is a geometric series and the multiplier is 1.03 per year but it is compounded monthly which makes it difficult.

Answer by stanbon(75887)

1000 euros is invested at 3% per annum interest, compounded monthly. Calculate the minimum number of months required for the value to exceed 1300 euros.
-------------
A(t) = P(1+(r/n))^(nt)
-----
1300 = 1000(1+(0.03/12))^(12*t)
-----
1.3 = (1.0025)^(12t)
----
12t*log(1.0025) = log(1.2)
----
12t = log(1.2)/log(1.0025)
12t = 73.02
# of months is 73.02
=========================
t would be the number of years
=========================
Cheers,
Stan H.
================